# TEDi Policy: Temporally Entangled Diffusion for Robotic Control

### v1

@@ 4.3 Push-T Evaluation
Here, we want to examine the closed-loop performance of TEDi Policy for the Push-T task and compare it to Diffusion Policy as a baseline. For Diffusion Policy, we keep the settings used by Chi et al. [1] by setting the prediction horizon $T_{p}$ to 16 and action horizon $T_{a}$ to $8$. We set the maximum number of epochs to 1000 since the performance typically converges before the end of training. This leads to slightly different results for Diffusion Policy than reported in the original paper. As two or more complete chunks are required in the buffer to speed up TEDi Policy prediction, we increase $T_{p}$ for our method. In the case of $T_{a}=8$, this requires a prediction horizon larger than $(T_{o}-1)+2\cdot 8=17$. Due to a restriction of the U-Net, we require the input to be divisible by 4 and therefore choose $T_{p}=20$. The results are reported in Table 1. We also measure the time taken to predict the action for each method.

### v2

@@ 4.3 Push-T Evaluation
Here, we want to examine the closed-loop performance of TEDi Policy for the Push-T task and compare it to Diffusion Policy as a baseline. For Diffusion Policy, we keep the settings used by Chi et al. [1] by setting the prediction horizon $T_{p}$ to 16 and action horizon $T_{a}$ to $8$. We set the maximum number of epochs to 1000 since the performance typically converges before the end of training. This leads to slightly different results for Diffusion Policy than reported in the original paper. As two or more complete chunks are required in the buffer to speed up TEDi Policy prediction, we increase $T_{p}$ for our method. In the case of $T_{a}=8$, this requires a prediction horizon larger than $(T_{o}-1)+2\cdot 8=17$. Due to a restriction of the U-Net, we require the input to be divisible by 4 and therefore choose $T_{p}=20$. The results are reported in Table 1. We also measure the time taken to predict an action chunk for each method.
@@ 5 Conclusion
Appendix A Sampling with incomplete chunks in the buffer
As shown by Chi et al. [1], values of $T_{o}$ and $T_{a}$ affect performance, and to do a fair comparison between the TEDi Policy and Diffusion Policy, we want to fix these parameters to $2$ and $8$, respectively. However, when fixing these parameters, there are no configurations possible if we respect both (i) the restriction that the input to the U-Net has to be divisible by 4 and (ii) we predict a multiple of $8$ actions into the future. Specifically, we would have to satisfy both
$T_{p}+(T_{o}-1)\equiv 0\pmod{T_{a}}\quad\text{and}\quad T_{p}\equiv 0\pmod{4}.$
Inserting $T_{o}=2$ and $T_{a}=8$, we get
$T_{p}+1\equiv 0\pmod{8}\quad\text{and}\quad T_{p}\equiv 0\pmod{4}$
which does not have any solutions. Any number divisible by $4$ is either $\equiv 0$ or $\equiv 4\bmod 8$. Therefore, we need to allow for values of $T_{p}-1$ that are not divisible by $T_{a}$, that is, we want to allow for incomplete chunks in the buffer. To allow for this, we modify Algorithm 2 slightly. The changes are highlighted in blue.
Algorithm 2 TEDi Policy Sampling
1:  if first step then
[…]
3:  end if
4:  while first chunk not clean do
6:  end while
9:  $\mathbf{B}\leftarrow\mathbf{B}[:-((T_{p}-(T_{o}-1))\text{\%}T_{a})]$
10:  $\mathbf{k}\leftarrow\mathbf{k}[:-((T_{p}-(T_{o}-1))\text{\%}T_{a})]$triangleright Remove the excess steps
11:  Sample $\mathbf{z}\sim\mathcal{N}(\mathbf{0}_{T_{a}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+((Tp-(To-1))\text{\%}Ta)}},\mathbf{I}_{T_{a}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+((Tp-(To-1))\text{\%}Ta)}})$
12:  $\mathbf{B}.\text{append}(\mathbf{z}),\mathbf{k}.\text{append}([N-1]_{\times T_{a}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+((Tp-(To-1))\text{\%}Ta)}})$triangleright Append $T_{a}$ steps of noise
13:  $\mathbf{k}_{\text{next}}\leftarrow\mathbf{k}[T_{a}]$triangleright Get diffusion level for next chunk
[…]
15:  return  $\mathbf{A}_{t}$